### 02 Mar 2026 v6

@@ 1 Introduction
On the other hand, for the $Eagle^{*}$ encryption algorithm, the key is divided into two independent parts, and the ciphertext can also be seen as two independent parts. One part of the ciphertext is completely determined by the intermediate random number which is completely unrelated to the plaintext. The other part of the ciphertext is determined by the intermediate random number together with the plaintext sequence, but it is completely indistinguishable whether the uncertainty of the ciphertext brought by the random number or brought by different plaintexts. This theoretically ensure that $Eagle^{*}$ can resist all forms of linear attacks and differential attacks.
@@ 2 Notations
$m*\delta$ - Excute the right multiplication of $\delta$ by $m$ introduced in chapter 5.2.
@@ 5.3.1 $Eagle^{*}$ key generator
The key is divided into two independent variables, denoted as $\delta_{1}$ and $\delta_{2}$, where $\delta_{1}$ and $\delta_{2}$ both have odd number of bits with 1.
Assuming $\delta_{1},\delta_{2}$ are all $L$ bits, the length of the entire key can be considered as 2L22L-2 bits, this is because each variable has a loss of 1 bit.
@@ 5.3.2 $Eagle^{*}$ encryption algorithm
We ramdomly generate $K$ groups number of $L$ bits, denoted as $R=(R_{1}$, $R_{2}$, …, $R_{K})$.
@@
0:Keys: $(\delta_{1},\delta_{2})$. Plaintext sequences: $P=(P_{1},P_{2},...,P_{K})$. Random number sequences: $R=(R_{1},R_{2},...,R_{K})$.
0:Ciphertext sequences: $C=(C_{1},C_{2},...,C_{K},S_{1},S_{2},...,S_{K})$.
1:for $i=1;i\leq K;i++$ do
2:  $X_{i}\leftarrow R_{i}*\delta_{1}*\delta_{1}$
3:  $C_{i}\leftarrow X_{i}\oplus P_{i}$
@@
5:for $j=1;j\leq K;j++$ do
6:  $S_{j}\leftarrow R_{j}*\delta_{2}*\delta_{2}$
7:end for
@@
$P_{1}$$R_{1}$$R_{1}$$X_{1}$$S_{1}$$C_{1}=P_{1}\oplus X_{1}$$P_{2}$$R_{2}$$R_{2}$$X_{2}$$S_{2}$$C_{2}=P_{2}\oplus X_{2}$$P_{K}$$R_{K}$$R_{K}$$X_{K}$$S_{K}$$C_{K}=P_{K}\oplus X_{K}$$\scriptstyle{R_{1}*\delta_{1}*\delta_{1}}$$\scriptstyle{R_{1}*\delta_{2}*\delta_{2}}$$\scriptstyle{R_{2}*\delta_{1}*\delta_{1}}$$\scriptstyle{R_{2}*\delta_{2}*\delta_{2}}$$\scriptstyle{R_{K}*\delta_{1}*\delta_{1}}$$\scriptstyle{R_{K}*\delta_{2}*\delta_{2}}$
As shown from the above figure, the ciphertext length is two times the plaintext length.
@@ 5.3.3 $Eagle^{*}$ decryption algorithm
For $Eagle^{*}$ decryptors, the ciphertext sequences are $C=($$C_{1}$, $C_{2}$, …, $C_{K}$, $S_{1}$, $S_{2}$, …, $S_{K})$. Here, the ciphertext sequences can be divided into two independent parts $(C_{1},C_{2}$, …, $C_{K})$ and $(S_{1},S_{2}$, …, $S_{K})$, first we use $(\delta_{2})$ and $(S_{1},S_{2},...,S_{K})$ to calculate the random number sequences $R=(R_{1},R_{2},...,R_{K})$, then we use the key $(\delta_{1})$ and the ciphertexts $(C_{1},C_{2},...,C_{K})$ and the random number sequences $R=(R_{1},R_{2},...,R_{K})$ to decrypt $P=(P_{1},P_{2},...,P_{K})$. The specific process is detailed in the following algorithm.
@@
0:Key: $(\delta_{1},\delta_{2})$. Ciphertext sequences: $C=(C_{1},C_{2},...,C_{K},S_{1},S_{2},...,S_{K})$.
@@
1:for $j=1;j\leq K;j++$ do
2:  $R_{j}\leftarrow S_{j}/\delta_{2}/\delta_{2}$
3:end for
4:for $i=1;i\leq K;i++$ do
5:  $X_{i}\leftarrow R_{i}*\delta_{1}*\delta_{1}$
6:  $P_{i}\leftarrow X_{i}\oplus C_{i}$
7:end for
@@ 5.3.4 Basic Security Analysis
Before delving into the security of the $Eagle^{*}$ encryption algorithm, we first introduce two sets of encryption algorithms, named A-Encrypt algorithm and B-Encrypt algorithm.
@@
0:Key: $(\delta_{2})$. Plaintext sequences: $R=(R_{1},R_{2},...,R_{K})$.
@@
1:for $j=1;j\leq K;j++$ do
2:  $S_{j}\leftarrow R_{j}*\delta_{2}*\delta_{2}$
@@
0:Key: $(\delta_{2})$. Ciphertext sequences: $C=(S_{1},S_{2},...,S_{K})$.
0:Plaintext sequences: $R=(R_{1},R_{2},...,R_{K})$.
1:for $j=1;j\leq K;j++$ do
2:  $R_{j}\leftarrow S_{j}/\delta_{2}/\delta_{2}$
3:end for
It is obvious that A-Encrypt, as an encryption algorithm, can correctly encrypt any plaintext and decrypt any ciphertext with any given key.
The encryption process of the B-Encrypt algorithm is as follows
Algorithm 10 B-Encrypt.
0:Key: $(\delta_{1})$. Plaintext sequences: $R=(R_{1},R_{2},...,R_{K})$.
0:Ciphertext sequences: $X=(X_{1},X_{2},...,X_{K})$.
1:for $i=1;i\leq K;i++$ do
2:  $X_{i}\leftarrow R_{1}*\delta_{1}*\delta_{1}$
3:end for
The decryption process of the B-Encrypt algorithm is as follows
Algorithm 11 B-Decrypt.
0:Key: $(\delta_{1})$. Ciphertext sequences: $X=(X_{1},X_{2},...,X_{K})$.
@@
[Theorem 11] In the $Eagle^{*}$ encryption algorithm, when know the plaintext $P=(P_{1},P_{2},...,P_{K})$ and the corresponding ciphertext $C=(C_{1},C_{2}$, …, $C_{K}$, $S_{1}$, $S_{2}$, …, $S_{K}$), the problem of cracking its key is equivalent to cracking their respective keys under the condition of knowing only the ciphertext $(S_{1},S_{2},...,S_{K})$ of A-Encrypt algorithm and only the ciphertext $(X_{1},X_{2},...,X_{K})$ of B-Encrypt algorithm.
@@
According to the definition of B-Encrypt, we have $X_{i}=P_{i}\oplus C_{i}$, know $P=(P_{1},P_{2},...,P_{K})$ and $C=(C_{1},C_{2},...,C_{K},S_{1},S_{2},...,S_{K})$ is equivalent to knowing $(X_{1},X_{2},...,X_{K})$ and $(S_{1},S_{2},...,S_{K})$. on the other hand, the parameters and keys used by A-Encrypt and B-Encrypt are completely independent, the conclusion of the theorem is obviously valid.
@@
According to the conclusion of Theorem 7, A-Encrypt and B-Encrypt obviously satisfies the above three conditions. This means that A-Encrypt and B-Encrypt can only use exhaustive search to crack the key when only knowing their ciphertexts, and there are no effective algorithms.
However, when putting A-Encrypt and B-Encrypt together, since they have a common plaintext sequences as $(R_{1},R_{2},...,R_{K})$, it is currently difficult to determine whether exhaustive search can only be used to crack the keys, this requires further analysis of the relationship between variables $(\delta_{1},\delta_{2})$ and the known plaintext-ciphertext pairs.
In fact, when either $\delta_{1}$ or $\delta_{2}$ is fixed, without loss of generality we fix $\delta_{1}$, $(R_{1},R_{2},...,R_{K})$ can be directly solved by $R_{i}=X_{i}/\delta_{1}/\delta_{1}$. To solve $\delta_{2}$, we need to solve the equation $R_{i}*\delta_{2}*\delta_{2}=S_{i}$, since this equation is a nonlinear equation about $\delta_{2}$.
As long as Assertion 1 holds, for the $Eagle^{*}$ algorithm, under the condition of knowing any plaintext-ciphertext, the attacker can only exhaustively search every possible value in the two sets of keys $(\delta_{1},\delta_{2})$, which satisfies the condition of practical security.
@@ 5.3.5 Linear attack analysis to $Eagle^{*}$ encryption algorithm
In the $Eagle^{*}$ encryption algorithm, the entire encryption system is divided into two independent parts, each with its own independent key and parameters. The ciphertext of the encryption system that includes the original plaintext is generated by mixing the original plaintext with random parameters, without knowing the random parameters, ciphertext can be regarded as completely independent of the original plaintext, which means that with random parameter changes, ciphertext can equally probability traverse any text in the ciphertext space. And the other part of the ciphertext is completely generated by random parameters, completely unrelated to the original plaintext. This can be expressed as

### 14 Mar 2026 v7

@@ 1 Introduction
On the other hand, for the $Eagle^{*}$ encryption algorithm, the key is divided into three independent parts, and the ciphertext can also be seen as three independent parts. Two parts of the ciphertext is completely determined by the intermediate random number which is completely unrelated to the plaintext. The third part of the ciphertext is determined by the intermediate random number together with the plaintext sequence, but it is completely indistinguishable whether the uncertainty of the ciphertext brought by the random number or brought by different plaintexts. This theoretically ensure that $Eagle^{*}$ can resist all forms of linear attacks and differential attacks.
@@ 2 Notations
$m*\delta$ - Excute the right multiplication of $m$ with $\delta$ introduced in chapter 5.2.
$m*\delta^{2}$ - Excute the right multiplication of $m$ with $\delta$ twice as $m*\delta*\delta$.
$s/\delta$ - Excute the right division of $s$ to $\delta$ introduced in chapter 5.2.
$s/\delta^{2}$ - Excute the right division of $s$ to $\delta$ twice as $s/\delta/\delta$.
@@ 5.2 The multiplication operation of binary
Without causing ambiguity, we denote the operation of performing two right multiplications by $\delta$ on $R$ as $R*\delta^{2}$.
Similarly, we can also introduce the division operator //, using the notation $S/\delta=R$ to represent $R*\delta=S$, and using $S/\delta^{2}$ to represent performing two division operations on $S$ with respect to $\delta$. The calculation for division is as follows.
Algorithm 6 calculation of $m/\delta$.
0:Input: $s,\delta$.
0:Output: $s/\delta$.
[…]
5:return $m$.
@@ 5.3.1 $Eagle^{*}$ key generator
The key is divided into three independent variables, denoted as $\delta_{1}$, $\delta_{2}$ and $\delta_{3}$, where $\delta_{1}$, $\delta_{2}$ and $\delta_{3}$ both have odd number of bits with 1.
Assuming $\delta_{1},\delta_{2},\delta_{3}$ are all $L$ bits, the length of the entire key can be considered as 3L33L-3 bits, this is because each variable has a loss of 1 bit.
@@ 5.3.2 $Eagle^{*}$ encryption algorithm
We randomly generate $2(K+3)$ groups number of $L$ bits, denoted as $x=(x_{1}$, $x_{2}$, …, $x_{K+3})$, $y=(y_{1}$, $y_{2}$, …, $y_{K+3})$, where $x_{i}$ and $y_{i}$ have odd number of bits different.
Then we randomly generate another four number as $P_{K+1},P_{K+2},P_{K+3}$ and $S_{0}$.
@@
0:Keys: $(\delta_{1},\delta_{2},\delta_{3})$. Plaintext sequences: $P=(P_{1},P_{2},...,P_{K})$. Random number sequences: $x=(x_{1},x_{2},...,x_{K+3})$,                $y=(y_{1},y_{2},...,y_{K+3})$.Random number: $P_{K+1},P_{K+2},P_{K+3}$ and $S_{0}$.
0:Ciphertext sequences: $C=(C_{1},C_{2},...,C_{K+3},S_{K+3}$                 $Sx_{1},Sx_{2},...,Sx_{K+3},$                 $Sy_{1},Sy_{2},...,Sy_{K+3})$
1:for $i=1;i\leq K+3;i++$ do
2:  $Sx_{i}\leftarrow x_{i}*\delta_{1}*\delta_{1}$
3:  $Sy_{i}\leftarrow y_{i}*\delta_{2}*\delta_{2}$
@@
5:for $j=1;j\leq K+3;j++$ do
6:  $(C_{j},S_{j})\leftarrow\mathbb{E}_{x_{j},y_{j}}(P_{j},S_{j-1})$
7:  if $j<K+3$ then
8:   $S_{j}\leftarrow S_{j}*\delta_{3}*\delta_{3}$
9:  end if
10:end for
@@
$S_{0}$$x_{1}*\delta_{1}^{2}=Sx_{1};y_{1}*\delta_{2}^{2}=Sy_{1}$$S_{1}*\delta_{3}^{2}$$C_{1}$$x_{2}*\delta_{1}^{2}=Sx_{2};y_{2}*\delta_{2}^{2}=Sy_{2}$$S_{2}*\delta_{3}^{2}$$C_{2}$$S_{K+2}*\delta_{3}^{2}$$C_{K+2}$$x_{K+3}*\delta_{1}^{2}=Sx_{K+3};y_{K+3}*\delta_{2}^{2}=Sy_{K+3}$$S_{K+3}$$C_{K+3}$$\scriptstyle{P_{1},x_{1},y_{1}}$$\scriptstyle{P_{2},x_{2},y_{2}}\scriptstyle{......}$$\scriptstyle{P_{K+3},x_{K+3},y_{K+3}}$
As shown from the above figure, the ciphertext length is about three times the plaintext length.
@@ 5.3.3 $Eagle^{*}$ decryption algorithm
For $Eagle^{*}$ decryptors, the ciphertext sequences are $C=($$C_{1}$, $C_{2}$, …, $C_{K+3}$, $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+3}$, $Sy_{1}$, $Sy_{2}$, …, $Sy_{K+3})$ and $S_{K+3}$. Here, the ciphertext sequences can be divided into three independent parts $(C_{1},C_{2}$, …, $C_{K+3},S_{K+3})$ and $(Sx_{1},Sx_{2}$, …, $Sx_{K+3})$ and $(Sy_{1},Sy_{2}$, …, $Sy_{K+3})$, First we use $(\delta_{1})$ and $(Sx_{1},Sx_{2},...,Sx_{K+3})$ to calculate the random number sequences $x=(x_{1}$, $x_{2}$, …, $x_{K+3})$, then we use $(\delta_{2})$ and $(Sy_{1}$, $Sy_{2}$, …, $Sy_{K+3})$ to calculate the random number sequences $y=(y_{1}$, $y_{2}$, …, $y_{K+3})$, then we use the key $(\delta_{3})$ and the ciphertexts $(C_{1}$, $C_{2}$, …, $C_{K+3}$, $S_{K+3})$ and the random number sequences $x=(x_{1}$, $x_{2}$, …, $x_{K+3})$ and $y=(y_{1}$, $y_{2}$, …, $y_{K+3})$ to decrypt $P=(P_{1}$, $P_{2}$, …, $P_{K})$. The specific process is detailed in the following algorithm.
@@
0:Key: $(\delta_{1},\delta_{2},\delta_{3})$. Ciphertext sequences: $C=(C_{1},C_{2},...,C_{K+3},S_{K+3},$                 $Sx_{1},Sx_{2},...,Sx_{K+3},$                 $Sy_{1},Sy_{2},...,Sy_{K+3})$.
@@
1:for $i=1;i\leq K+3;i++$ do
2:  $x_{i}\leftarrow Sx_{i}/\delta_{1}/\delta_{1}$
3:  $y_{i}\leftarrow Sy_{i}/\delta_{2}/\delta_{2}$
4:end for
5:for $j=K+3;j\geq 1;j$-- do
6:  $(P_{j},S_{j-1})\leftarrow\mathbb{D}_{x_{j},y_{j}}(C_{j},S_{j})$
7:  $S_{j-1}\leftarrow S_{j-1}/\delta_{3}/\delta_{3}$
8:end for
@@ 5.3.4 Basic Security Analysis
Before delving into the security of the $Eagle^{*}$ encryption algorithm, we first introduce another temporary encryption algorithms, named A-Encrypt algorithm.
@@
0:Key: $(\delta)$. Plaintext sequences: $R=(R_{1},R_{2},...,R_{K})$.
@@
1:for $i=1;i\leq K;i++$ do
2:  $S_{i}\leftarrow R_{i}*\delta*\delta$
@@
0:Key: $(\delta)$. Ciphertext sequences: $C=(S_{1},S_{2},...,S_{K})$.
@@
[Theorem 11] In the $Eagle^{*}$ encryption algorithm, when know the plaintext $P=(P_{1},P_{2},...,P_{K})$ and the corresponding ciphertext $C=(C_{1},C_{2}$, …, $C_{K+3}$, $S_{K+3}$, $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+3}$, $Sy_{1}$, $Sy_{2}$, …, $Sy_{K+3}$), the problem of cracking its key $(\delta_{1},\delta_{2},\delta_{3})$ is equivalent to cracking two respective keys $\delta_{1}$ and $\delta_{2}$ under the condition of knowing only the ciphertext $(Sx_{1},Sx_{2},...,Sx_{K+3})$ of A-Encrypt algorithm and only the ciphertext $(Sy_{1},Sy_{2},...,Sy_{K+3})$ of another A-Encrypt algorithm.
@@
According to the definition of $Eagle^{*}$, when know the plaintext $P=(P_{1},P_{2},...,P_{K})$ and the corresponding ciphertext $C=(C_{1},C_{2}$, …, $C_{K+3}$, $S_{K+3}$, $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+3}$, $Sy_{1}$, $Sy_{2}$, …, $Sy_{K+3}$), the problem of cracking its key $(\delta_{1},\delta_{2},\delta_{3})$ is equivalent to solve the following system of equations.
[…]
Where $x_{1},x_{2},...,x_{K+3},y_{1},y_{2},...,y_{K+3},P_{K+1},P_{K+2},P_{K+3},\delta_{1},\delta_{2}$ are unknown variables to be solved. $P_{1}$, $P_{2}$, …, $P_{K+3}$, $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+3}$, $Sy_{1}$, $Sy_{2}$, …, $Sy_{K+3}$, $S_{K+3}$ are known variables.
According to the conclusion of Theorem 3, each equation in the system of equations (5.5) can be regarded as an identity equation.
The unknown variables $x_{1}$, $x_{2}$, …, $x_{K+3}$ in the system of equations (5.6) can be equivalently regarded as the plaintexts of the A-encryption algorithm, with the known variables $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+3}$ being the ciphertexts of the A-encryption algorithm, and $\delta_{1}$ being the key. Solving the system of equations (5.6) is equivalent to crack the key $\delta_{1}$ under the condition of knowing only the ciphertext $(Sx_{1},Sx_{2},...,Sx_{K+3})$ of A-Encrypt algorithm.
Similarly, solving the system of equations (5.7) is equivalent to crack the key $\delta_{2}$ under the condition of knowing only the ciphertext $(Sy_{1},Sy_{2},...,Sy_{K+3})$ of another A-Encrypt algorithm.
@@
According to the conclusion of Theorem 7, A-Encrypt obviously satisfies the above three conditions. This means that A-Encrypt can only use exhaustive search to crack the key when only knowing their ciphertexts, and there are no effective algorithms.
Next, consider several possible exhaustive attack methods.
[Case 1] Only enumerate either $\delta_{1}$ or $\delta_{2}$.
It suffices to consider only the case where $\delta_{1}$ is exhausted, $x_{1}$, $x_{2}$, …, $x_{K+3}$ can be directly calculated from the system of equations (5.6), but this does not change the conclusion that each equation in the system of equations (5.5) is an identity equation.
[Case 2] Only enumerate $\delta_{3}$.
When only exhaust $\delta_{3}$, since both $x_{i}$ and $y_{i}$ are unknown, it is impossible to determine $S_{i}$ solely based on $S_{K+3}$, which brings no benefit to solving the system of equations (5.5), (5.6), and (5.7).
[Case 3] Only enumerate $(\delta_{1},\delta_{2})$.
When enumerating over $(\delta_{1},\delta_{2})$, we can directly calcute $x_{1}$, $x_{2}$, …, $x_{K+3}$ from the system of equations (5.6), and directly calculate $y_{1}$, $y_{2}$, …, $y_{K+3}$ from the system of equations (5.7). Substituting these calculated $x_{i}$ and $y_{i}$ into the system of equations (5.5), we can sequentially calculate the hidden variables $S_{1}$, $S_{2}$, …, $S_{K+2}$. Then we can solve the equation $S_{i}*\delta_{3}*\delta_{3}=S_{i+1}$ to obtain $\delta_{3}$, this equation is a nonlinear equation about $\delta_{3}$. As long as Assertion 1 holds, the only way to solve this equation is exhaustively enumerating $\delta_{3}$.
[Case 4] Only enumerate $(\delta_{1},\delta_{3})$ or $(\delta_{2},\delta_{3})$.
It suffices to consider only the case where $(\delta_{1},\delta_{3})$ is exhausted, $x_{1}$, $x_{2}$, …, $x_{K+3}$ can be directly calculated from the system of equations (5.6). Since $y_{i}$ is unknown, it is impossible to determine $S_{i}$ solely based on $x_{i}$ and $S_{K+3}$, which brings no benefit to solving the system of equations (5.5), (5.6), and (5.7).
From the above analysis, it can be seen that for any known plaintext-ciphertext correspondence, the correct key can only be found by exhaustively enumerating all $(\delta_{1},\delta_{2},\delta_{3})$.
Next, we need to consider another possible attack method. The attacker finds the relationship between the ciphertext to be decrypted and the ciphertexts- plaintexts in the sample from a large number of plaintext-ciphertext pairs, and directly decrypts or partially decrypts without cracking the key. The premise for this attack method to be effective is that there is some statistical relationship or some specific linear relationship between the plaintext and the ciphertext.
Due to the complete randomness of $x_{i}$ and $y_{i}$, there is no statistical relationship or any form of mathematical relationship between the plaintext $P_{i}$ and $Sx_{i}$. Similarly, there is no statistical relationship between $P_{i}$ and $Sy_{i}$. Furthermore, according to Theorem 3, $P_{i}$ and $C_{i}$ are also completely unrelated. In fact, if we disregard $S_{i}$ and $\delta_{3}$ (in reality, only $S_{K+3}$ and $\delta_{3}$ are known, all $S_{i}$ are invisible without solving for $x_{i}$ and $y_{i}$), the encryption process of $Eagle^{*}$ is equivalent to one-time pad. An attacker cannot find any patterns between plaintexts and ciphertexts from a large sample of plaintext-ciphertext pairs.
Lastly, it should be noted that during encryption, why do we need to randomly generate three additional sets of plaintexts $(P_{K+1},P_{K+2},P_{K+3})$ after $P_{K}$? This is because there exists an implicit relationship between $S_{K+3}$ and $P_{K+3}$ in the ciphertext, which is $S_{K+3}=P_{K+3}*(x_{K+3}\oplus y_{K+3})\oplus x_{K+3}^{+(L-1)}$. In order to make this implicit relationship invisible, we generated several additional sets of plaintext sequences when designing $Eagle^{*}$. The reason why we chose three sets is that there are three variables in the key, so we opted to generate three additional sets of plaintexts.
@@ 5.3.5 Linear attack analysis to $Eagle^{*}$ encryption algorithm
In the $Eagle^{*}$ encryption algorithm, the entire encryption system is divided into three independent parts, each with its own independent key and parameters. The ciphertext of the encryption system that includes the original plaintext is generated by mixing the original plaintext with random parameters, without knowing the random parameters, ciphertext can be regarded as completely independent of the original plaintext, which means that with random parameter changes, ciphertext can equally probability traverse any text in the ciphertext space. And the two other parts of the ciphertext is completely generated by random parameters, completely unrelated to the original plaintext. This can be expressed as